OpenStudy (anonymous):

how to simplify the square root of 50 step by step

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{25*2}=5\sqrt{2}\]

OpenStudy (anonymous):

Using calculator sqrt [50] = 7.07 Otherwise, sqrt [50] = sqrt [2 x 25] = sqrt [2] x sqrt [25] = sqrt [2] x 5 = 5 sqrt (2)

OpenStudy (anonymous):

what about 99

OpenStudy (anonymous):

\[\sqrt{9*11}\]

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

\[\sqrt{99}=\sqrt{9*11}=\]

OpenStudy (anonymous):

\[\sqrt{9}=3 \] SO \[3\sqrt{11}\]

OpenStudy (anonymous):